OpenStudy (anonymous):

Find the domain: I think factoring may be involved in this... \[g(x)=\frac{ 3 }{ x ^{2}-2x-15 }\]

OpenStudy (campbell_st):

well you need to factor the denominator and find the values of x that make the denominator zero. the values of x will be the restrictions in the domain. the domain is all real x, except the restrictions you find. hope that helps

OpenStudy (anonymous):

Before I can find the restrictions, is it mandatory that I factor the denominator first?

OpenStudy (campbell_st):

well the restrictions occur in the denominator... so to find them factoring your quadratic is an easy solution... of if that's a problem you can use the general quadratic formula

OpenStudy (campbell_st):

the factored form is \[x^2 - 2x -15 = (x - 5)(x +3)\] so find the values of x that make the equation equal zero... they will be the restrictions

OpenStudy (anonymous):

If I remember doing this correctly, then this factored would be: \(\ \dfrac{3}{x^2 -5x + 3x - 15} \) \(\ \dfrac{3}{x(x - 5) + 3(x - 5)} \) Then, \(\ \dfrac{3}{(x - 5)(x + 3)}\) The restrictions are; x = 5, x = -3 \(\ \dfrac{3}{5^2 - 2(5) - 15} \implies \dfrac{3}{25 - 10 - 15}\implies \dfrac{3}{0} \)

OpenStudy (anonymous):

oh my goodness you guys ares geniuses e.e thank you...I think I'll try the next problem myself